Riverhawk Sales, Inc. had a net income of $15,000 on sales of $100,000 last year. Its balance sheet shows debt of $20,000 and equity of $72,000. If the total dividend paid was $8,000, what is its sustainable growth rate
Answer #1
Return on equity (ROE) = Net income / Equity = 15000 / 72000 20.83%
Retention ratio (RR) = ( Net income - Dividend ) / Net income = ( 15000 - 8000 ) / 15000 46.67%
Sustainable growth rate = ROE * RR = 20.83% * 46.67% 9.72%